OpenStudy (anonymous):

\[\text{erf}(\infty)=1\]

hero (hero):

Huh?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

lol...sorry...i'll fix it now

hartnn (hartnn):

and erf(-infinity)=-1 or u wanted to prove that?

OpenStudy (anonymous):

just for remembring as one of my friends asked

OpenStudy (unklerhaukus):

\[\text{erf}(\infty)=\frac{2}{\sqrt \pi}\int\limits_0^\infty e^{-u^2}\text du=1\]

hartnn (hartnn):

@UnkleRhaukus ,which graph is that? its not of erf....

OpenStudy (unklerhaukus):